one side of a triangle is 4cm longer and another 3cm shorter than the third side.if the perimeter of the triangle is 43cm, find the length of the three sides.
Answer by mananth(16946)

let third side =x
one side = x+4
second side=x-3
...
x+x+4+x-3 =43
3x+1=43
3x=42
/3
x=14 cm third side
one side = 17cm
second side = 11 cm
...
m.ananth@hotmail.ca